LBC Tech Chip IA Commissioning - 05 May 2010 UT

Night 1

Observers: M.Pedani (LBTO)
Telescope Operator: SAllanson (LBTO), JHill (LBTO)
Telescope Support: DAshby(LBTO), MMidkiff(LBTO)
Instrument Support: DOfficer (LBTO)
LBC team: RSpeziali (INAF), ABaruffolo(INAF)

Summary

Photometric Sky. T=8.1 at beginning, Wind up to 9m/s. At the end of the night T=6.5C
IA tonight is -371.9, pretty similar to Last Night Value with LUCI (it was -398)
DOFPIA now has 250nm threshold for convergence with the exception of Z7, Z8 (400nm).
Despite the conditions were stable, it always took 7-9 iterations before converging both at Blue and Red!

Filters Focus Offsets obtained tonight:
Note that on Apr10 the first LBCB Filters Focus Offsets were measured after correcting the Z22 zeropoint.

LBCB:

U-Bessel = 0.344

U-Spec = 0.443404
U-Spec = 0.447752

U-Spec Average = 0.446

V-Bessel = -0.0118884
V-Bessel = -0.0186031
V-Bessel = 0.0079274
V-Bessel = 8.60074E-4

V_Bessel Average = 0.000

B-Bessel = 0.0342378
B-Bessel = 0.040276
B-Bessel = 0.0678034
B-Bessel = 0.0428277

B-Bessel Average = 0.046

g-sdss = 0.065

r-sdss = 0.091

LBCR:

z-sdss = -0.023

R-Bessel = 0.023192
R-Bessel = -0.0073283
R-Bessel = 0.0241658

R-Bessel Average = 0.013

I-Bessel = -0.047

i-sdss = -0.021

V-Bessel = 0.079

r-sdss = 0.046

Y-Fan = -0.023513
Y-Fan = -0.0277874

Y_Fan Average = -0.026

F972 = 0.13904
F972 = 0.138014

F972 Average = 0.139

Open

02:35 Open, Photometric, T_out = 8.1C Humidity = 13% T_mirror=6.7C Wind=9 m/s

Details

03:00 Problem with Left Mirror Cell Actuators Voltage

03:20 Problem with LBC HK. Lbckill and LbcStart fixed it!

03:35 WT_290 and DOFPIA. AZ dropped to Velocity Mode. Not tracking ?

03:40 ACT_290 and DOFPIA again. EL = 79 T_mirror=7.2C

03:55 We are collimated! Seeing Estimated at Red = 0.6", Seeing estimated at Blue = 1.7"

03:57 WT10_307 Checking IA at High Elevation; CA=-48.7 IE=-40.2

04:00 ACT0086 to check IA at Low Elevation (21 Deg.) ; Hit DX Mirror Limits at Z. IA=-371.9 CA=-48.7 IE=-37.4

04:10 ACT309 and DOFPIA to Collimate at EL=+80;

04:30 ACT309 and Start Testing Tech Chip IA

04:35 ACT309 Start 300s exposure Blue+Red; lbcb.20100505.043620 and lbcr.20100505.043617
John sending Z4 Corrections by hand to see how the Tech Chip Image does change

04:37:04 Sent +400nm Z4 at Blue and Red
04:38:05 Sent -800nm Z4 at Blue and Red; not applied at Red ?
04:39:20 Sent +400nm Z4 at Blue and Red
04:40:23 Sent +800nm Z4 at Blue and Red
04:41:15 Sent -1600nm Z4 at Blue and Red; not applied at Red?

04:43 ACT309 EL=+85; DOFPIA again

04:55 T_out = 7.8 T_Mirror = 8.0C Wind ~11m/s

04:57 ACT309 Start 300s exposure Blue+Red; lbcb.20100505.045746 and lbcr.20100505.045743

04:58:50 Sent +400nm Z4 at Blue and Red
04:59:38 Sent -800nm Z4 at Blue and Red
05:00:18 Sent +400nm Z4 at Blue and Red
05:01:02 Sent +800nm Z4 at Blue and Red
05:01:47 Sent -1600nm Z4 at Blue and Red
05:02:25 Sent +800nm Z4 at Blue and Red

05:04:55 LBCKill and LBCStart; John changes the AGuide Minimum Correction to be sent at Red to +5"
From now on no correction will be sent from LBCR!

05:08 ACT309 and DOFPIA at EL=+86

05:30 ACT309 offsetted by -140"RA Start 300s exposure Blue+Red; lbcb.20100505.053107 and lbcr.20100505.053103
we should have a brigh star in tech Chip now!
PA = +9.6 why ?

05:37 ACT309 offsetted by -140"RA Start 300s exposure Blue+Red; lbcb.20100505.053647 and lbcr.20100505.053647

05:44 ACT309 Offsetted by +790" to put the Pointing Star into the Tech Chips!
Some tests to realize that a -860" offset at RA is needed to put the center of the field into the tech Chip

05:58 ACT309 to adjust Pointing; CA= -60 IE=-58

06:02 ACT309 Offsetted by -860" to have Point Star on tech chip; almost there it could be -890:

06:11 LBCR in weird status does not respond; lbckill and lbsstart

Got RPC Error; Updated Issue Track #2397

06:27 ACT309 and DOFPIA again!

06:31 T_out = 7.2C T_Mirror=7.7C wind up to 10m/s

06:38 ACT309 offsetted RA 600s exposure; Blue+Red; lbcb.20100505.063916 and lbcr.20100505.063913

06:40:25 Sent +400nm Z4 at Blue and Red
06:41:22 Sent -800nm Z4 at Blue and Red
06:42:22 Sent +400nm Z4 at Blue and Red
06:43:22 Sent +800nm Z4 at Blue and Red
06:44:22 Sent -1600nm Z4 at Blue and Red
06:45:22 Sent +800nm Z4 at Blue and Red
06:46:21 Sent +1600nm Z4 at Blue and Red
06:47:21 Sent -3200nm Z4 at Blue and Red

06:51 Slewing to WT10_311 to recollimate at EL=+84; DOFPIA

07:16 RB_RV_superfoc while Baru is working on Tech Chip Images

Blue-V : Average best focus of -0.0118884 with FWHM of 2.52
Red-R : Average best focus of 3.39140E-4 with FWHM of 2.52

07:30 Again WT10_311 and DOFPIA

07:40 RB_zB_superfoc

Blue-B: Average best focus of 0.0342378 with FWHM of 2.93
Red-z : parabola too shifted toward negative values..

07:56 RB_zB_superfoc again to check Red; edited Starting point for Red to -0.19mm instead from -0.09mm

Don't see any change in the z-sdss parabola !!

08:12 RB_zB_Coarse_superfoc

Blue-B: Average best focus of 0.040276 with FWHM of 2.53
Red-z: Average best focus of -0.0234237 with FWHM of 2.27

08:20 Slew to WT10_313 El=+86

08:28 RB_RV_superfoc

Blue-V : Average best focus of -0.0186031 with FWHM of 3.40
Red-R : Average best focus of 0.023192 with FWHM of 3.24

08:36 DOFPIA at El=+83

08:42 RB_Ir_superfoc

Blue-r: Average best focus of 0.0912272 with FWHM of 3.37
Red-I : Average best focus of -0.0470204 with FWHM of 2.57

08:49 DOFPIA at El=+80

08:53 RB_ig_superfoc

Blue-g: Average best focus of 0.0650294 with FWHM of 3.05
Red-i : Average best focus of -0.0211973 with FWHM of 2.27

09:00 DOFPIA at El=+78

09:07 RB_VV_superfoc

Blue-V: Average best focus of 0.0079274 with FWHM of 2.80
Red-V: Average best focus of 0.0791554 with FWHM of 3.93

09:15 DOFPIA at El=+75

09:19 RB_rV_superfoc

Blue-V: Average best focus of 8.60074E-4 with FWHM of 2.76
Red-r: Average best focus of 0.0457015 with FWHM of 2.61

09:26 DOFPIA at EL=+72

09:31 RB_YU_superfoc

Blue-U: Does not converge! Look close to 0.275
Red-Y: Average best focus of -0.023513 with FWHM of 2.95

09:40 DOFPIA at El=+70

09:45 RB_YU_superfoc

Blue-U: Average best focus of 0.343611 with FWHM of 3.59
Red-Y: Average best focus of -0.0277874 with FWHM of 2.21

09:50 DOFPIA at El=+66

09:51 RB_RUs_superfoc

Blue-Us: Average best focus of 0.443404 with FWHM of 3.33
Red-R: Average best focus of -0.0073283 with FWHM of 3.09

10:15 Slew to ACT300 El+84 and DOFPIA

10:23 RB_RUs_superfoc

Blue-Us: Average best focus of 0.447752 with FWHM of 3.08
Red-R: Average best focus of 0.0241658 with FWHM of 2.80

10:32 ACT300 DOFPIA at El=+81

10:40 RB_F972B_superfoc

Blue-B: Average best focus of 0.0678034 with FWHM of 3.42
Red-F972: Average best focus of 0.13904 with FWHM of 3.08

10:50 ACT300 DOFPIA at EL=+78

10:55 RB_F972B_superfoc

Blue-B: Average best focus of 0.0428277 with FWHM of 2.46
Red-F972: Average best focus of 0.138014 with FWHM of 2.70

11:05 DOFPIA and centering a suitable star into the Blue Tech Chip

11:15 Test Guide 600s; ACT300 Offsetted to get one star into the Blue Tech Chip

11:17:01 Sent +400nm Z4 at Blue and Red
11:18:01 Sent -800nm Z4 at Blue and Red
11:19:02 Sent +400nm Z4 at Blue and Red
11:20:00 Sent +800nm Z4 at Blue and Red
11:20:55 Sent -1600nm Z4 at Blue and Red
11:21:44 Sent +800nm Z4 at Blue and Red
11:22:37 Sent +1600nm Z4 at Blue and Red
11:23:33 Sent -3200nm Z4 at Blue and Red
11:24:29 Sent +1600nm Z4 at Blue and Red

11:30 Slew to B3.2 2120+0006 and DOFPIA at El=+45
11:41 Too Bright we close

-- MarcoPedani - 05 May 2010
Topic revision: r4 - 06 May 2010, MarcoPedani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback